The probability density function (pdf) of the normal distribution is given by 1 V2πσ2 fx (x)= 0 (-(x - 1)
1
Expert's answer
2021-09-06T16:23:16-0400

"f_x(x; \sigma^2)= \frac{1}{\sqrt{2 \pi \sigma^2}} e^{-\frac{x^2}{2 \sigma^2}}"

By factorization theorem , we can write it as

"=g(\sum_i^n x_i^2 \sigma^2)*h(x)"

Therefore

"T=\sum_1^n x_i^2" is sufficient statistic for "\sigma^2" where "h(x)=1"